[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Jul 13, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Jul 13 19:50:00 UTC 2016


HSU Shift Log for Wed Jul 13 2016 15:49:55 GMT-0400 (EDT)

Shift: H

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS/AWIPS2 issues:

Other Techincal issues: Noticed at 1800Z that the text TWOEP hadn't updated  
on the NHC website. Let Dave Z. know. Probably due to lingering problems at  
the TOC.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160713/d900e61b/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list